OpenStudy (jmartinez638):

Describe the continuity of the function f(x) = (x^2-16)/(x^2-9x+20). For each point of discontinuity, describe the graph's behavior in terms of asymptotes, removable discontinuities, etc.

OpenStudy (jmartinez638):

\[f(x) = \frac{x^{2}-16 }{x^{2}-9x+20}\]

OpenStudy (anonymous):

factor both the numerator and denominator. Anything that cancels out is a hole (removable discontinuity). From what's left after canceling, set the denominator equal to 0 to find the vertical asymptotes (infinite discontinuity)

OpenStudy (jmartinez638):

Ok, I'll do that. Thanks!

OpenStudy (jmartinez638):

Removable discontinuity at x-4, and vertical asymptote at x+5

OpenStudy (anonymous):

x - 4 = 0 x = 4 so removable disc. @ x = 4 x - 5 = 0 x = 5 vertical asymptote is x = 5